A sample containing Cd2+ and Mn2+ was treated with 60.1 ml of 0.0400M EDTA. Titration of the excess unreacted EDTA required 16.8ml of 0.013M Ca2+. The Cd2+ was displaced from EDTA by the addition of an excess of CN-. Titration of the newly freed EDTA required 10.4ml of 0.013M Ca2=. What are the concentrations of Cd2+ and Mn2+ in the original solution?
Answer #1

EDTA reactions with metal ions are 1:1 :

Cd2+ + EDTA <==> (Cd-EDTA)

Mn2+ + EDTA <==> (Mn-EDTA)

Ca2+ + EDTA <==> (Ca-EDTA)

This experiment involves back-titration as masking.

Sample containing Cd2+ and Mn2+ was treated with 60.1 ml of 0.0400M EDTA.

moles of EDTA =( 0.0400 mol/L*60.1 ml) / 1000 ml/L = 2.404*10-3 mole

Titration of the excess unreacted EDTA required 16.8ml of 0.013M Ca2+.

So, moles of Ca2+ of reacted = ( 0.013 mol/L* 16.8 ml) / 1000 ml/L =2.184*10-4 mole

so moles of ( Cd2+ +Mn2+ ) = 2.1856*10-3 mole

addition of excess of CN- , displaced Cd2+ from Cd-EDTA complex and liberated EDTA, free EDTA was back titrated with   Ca2+ :

moles of Ca2+ reacted = moles EDTA liberated = moles of Cd2+ present

=  ( 0.013 mol/L* 10.4 ml) / 1000 ml/L =1.352*10-4 mole

so, moles of Cd2+ present in initial solution : 1.352*10-4 moles

and moles of Mn2+ present in initial solution : = 2.1856*10-3 mole​​​​​​​ - 1.352*10-4 mole​​​​​​​

= 2.0504*10-3 moles

(you can calculate respective concentration if you volume of sample initially reacted )